# Strichartz estimates for the generalized Zakharov-Kuznetsov equation on ×$\mathbb{R}\times\mathbb{T}$ and applications

### v1

@@ 4.2. Improved LWP for mZK
(ii.3.3.1) $\lvert\xi_{i}+\xi_{j}\rvert\gtrsim\lvert\xi_{\text{max}}\rvert$, for all $(i,j)\in\{(1,2),(1,3),(2,3)\}$:
In this case, we have, due to the general assumption (ii.3), that
@@
which allows us to control up to $\frac{3}{2}-$ derivatives by means of the quantities $\langle\sigma_{i}\rangle$. Proceeding exactly as in case (ii.3.2.2), we thus arrive at (42), and this holds for every $s>\frac{1}{3}$, provided we choose $\varepsilon>0$ sufficiently small.
(ii.3.3.2) $\lvert\xi_{i}+\xi_{j}\rvert\gtrsim\lvert\xi_{\text{max}}\rvert$, for exactly two $(i,j)\in\{(1,2),(1,3),(2,3)\}$:
We may assume without loss of generality that $\lvert\xi_{1}+\xi_{2}\rvert\ll\lvert\xi_{\text{max}}\rvert$, and this together with (ii.3.3) immediately implies $\lvert\xi_{0}\rvert\sim\lvert\xi_{1}\rvert\sim\lvert\xi_{2}\rvert\sim\lvert\xi_{3}\rvert$. Moreover, $\lvert\xi_{0}\rvert\lesssim\lvert\xi_{0}+\xi_{1}\rvert$, $\lvert\xi_{0}\rvert\lesssim\lvert\xi_{2}+\xi_{3}\rvert$, $\lvert\xi_{0}\rvert\lesssim\lvert\xi_{0}+\xi_{2}\rvert$, $\lvert\xi_{0}\rvert\lesssim\lvert\xi_{1}+\xi_{3}\rvert$, and $\lvert\xi_{0}\rvert\lesssim\lvert\xi_{0}+\xi_{3}\rvert$ hold, and if the bilinear refinement is applicable in at least one of these configurations, we can proceed as in case (ii.3.1) to obtain the desired estimate (42) for all $s>\frac{3}{8}$. We are therefore left to examine the situation in which
$\begin{split}\lvert 3(\xi_{0}+\xi_{1})^{2}-(q_{0}+q_{1})^{2}\rvert&\ll\lvert\xi_{0}+\xi_{1}\rvert\sim\lvert\xi_{0}\rvert,\\ \lvert 3(\xi_{2}+\xi_{3})^{2}-(q_{2}+q_{3})^{2}\rvert&\ll\lvert\xi_{2}+\xi_{3}\rvert\sim\lvert\xi_{0}\rvert,\\ \lvert 3(\xi_{0}+\xi_{2})^{2}-(q_{0}+q_{2})^{2}\rvert&\ll\lvert\xi_{0}+\xi_{2}\rvert\sim\lvert\xi_{0}\rvert,\\ \lvert 3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}\rvert&\ll\lvert\xi_{1}+\xi_{3}\rvert\sim\lvert\xi_{0}\rvert,\\ \lvert 3(\xi_{0}+\xi_{3})^{2}-(q_{0}+q_{3})^{2}\rvert&\ll\lvert\xi_{0}+\xi_{3}\rvert\sim\lvert\xi_{0}\rvert\end{split}$ (45)
hold. We first observe: from (45) it follows that $\lvert\xi_{0}\rvert\sim\lvert\xi_{1}\rvert\sim\lvert\xi_{2}\rvert\sim\lvert\xi_{3}\rvert\sim\lvert(\xi_{1},q_{1})\rvert$, since if $\lvert\xi_{1}\rvert\ll\lvert(\xi_{1},q_{1})\rvert$, then from (ii.3) it would follow that $\lvert q_{0}\rvert\sim\lvert q_{1}\rvert\sim\lvert q_{2}\rvert\sim\lvert q_{3}\rvert\sim\lvert(\xi_{1},q_{1})\rvert$, which would contradict at least one of the inequalities in (45). Subtracting the first inequality in (45) from the second, and the third from the fourth, we obtain
$\lvert 3\xi_{0}\xi_{1}-q_{0}q_{1}\rvert\ll\lvert\xi_{0}\rvert$ (46)
@@
$\lvert 3\xi_{0}\xi_{2}-q_{0}q_{2}\rvert\ll\lvert\xi_{0}\rvert,$ (47)
which immediately leads us to $\lvert q_{0}\rvert\sim\lvert q_{1}\rvert\sim\lvert q_{2}\rvert\sim\lvert(\xi_{1},q_{1})\rvert$. Adding (46) and (47), we further deduce
$\lvert 3\xi_{0}(\xi_{1}+\xi_{2})-q_{0}(q_{1}+q_{2})\rvert\ll\lvert\xi_{0}\rvert,$ (48)
from which it follows that $\lvert q_{1}+q_{2}\rvert\ll\lvert\xi_{0}\rvert$, so that we can conclude $\lvert q_{3}\rvert\sim\lvert(\xi_{1},q_{1})\rvert$ as well. Altogether, we have thus confirmed that
$\lvert\xi_{0}\rvert\sim\lvert\xi_{1}\rvert\sim\lvert\xi_{2}\rvert\sim\lvert\xi_{3}\rvert\sim\lvert q_{0}\rvert\sim\lvert q_{1}\rvert\sim\lvert q_{2}\rvert\sim\lvert q_{3}\rvert\sim\lvert(\xi_{1},q_{1})\rvert.$ (49)
We now turn to an analysis of the signs. From (46), (47), (49), $\lvert\xi_{1}+\xi_{2}\rvert\ll\lvert\xi_{3}\rvert\sim\lvert\xi_{0}\rvert$, and $\lvert q_{1}+q_{2}\rvert\ll\lvert\xi_{0}\rvert$, we obtain the following possible sign configurations, all of which can be treated analogously.
$(\xi_{0},\xi_{1},\xi_{2},\xi_{3})$ $(q_{0},q_{1},q_{2},q_{3})$
In what follows, we shall restrict our attention to the case $(\xi_{0},\xi_{1},\xi_{2},\xi_{3})=(q_{0},q_{1},q_{2},q_{3})=(+,+,-,+)$. Incorporating the sign configuration into the inequalities established in (45) leads us to
$\lvert\sqrt{3}(\xi_{0}+\xi_{1})-(q_{0}+q_{1})\rvert\ll 1,$
$\lvert\sqrt{3}(\xi_{0}+\xi_{3})-(q_{0}+q_{3})\rvert\ll 1,$
and
$\lvert\sqrt{3}(\xi_{1}+\xi_{3})-(q_{1}+q_{3})\rvert\ll 1,$
and subsequent addition and subtraction of these inequalities gives
$\begin{split}\lvert\sqrt{3}\xi_{0}-q_{0}\rvert&\ll 1,\\ \lvert\sqrt{3}\xi_{1}-q_{1}\rvert&\ll 1,\\ \lvert\sqrt{3}\xi_{2}-q_{2}\rvert&\ll 1,\\ \lvert\sqrt{3}\xi_{3}-q_{3}\rvert&\ll 1.\end{split}$ (50)
With these estimates in mind, we now rewrite the resonance function as follows:
$R$ $=\xi_{0}(\xi_{0}^{2}+q_{0}^{2})-\xi_{1}(\xi_{1}^{2}+q_{1}^{2})-\xi_{2}(\xi_{2}^{2}+q_{2}^{2})-\xi_{3}(\xi_{3}^{2}+q_{3}^{2})$
$=4\xi_{0}^{3}-4\xi_{1}^{3}-4\xi_{2}^{3}-4\xi_{3}^{3}+\xi_{0}(-3\xi_{0}^{2}+q_{0}^{2})-\sum_{i=1}^{3}\xi_{i}(-3\xi_{i}^{2}+q_{i}^{2})$
$=12(\xi_{1}+\xi_{2})(\xi_{1}+\xi_{3})(\xi_{2}+\xi_{3})+\xi_{0}(-3\xi_{0}^{2}+q_{0}^{2})-\sum_{i=1}^{3}\xi_{i}(-3\xi_{i}^{2}+q_{i}^{2}).$
It now follows that
$\lvert R\rvert\geq 12\lvert\xi_{1}+\xi_{2}\rvert\lvert\xi_{1}+\xi_{3}\rvert\lvert\xi_{2}+\xi_{3}\rvert-6\lvert\xi_{\text{max}}\rvert\sum_{i=0}^{3}\lvert\sqrt{3}\xi_{i}-q_{i}\rvert\lvert\sqrt{3}\xi_{i}+q_{i}\rvert,$ (51)

### v3

@@ 4.2. Improved LWP for mZK
(ii.3.3.1) ${\lvert}\xi_{\text{max}}{\rvert}\ll{\lvert}(\xi_{1},q_{1}){\rvert}$:
In this situation, it follows from the active constraint (ii.3) that
${\lvert}q_{1}{\rvert}\sim{\lvert}q_{2}{\rvert}\sim{\lvert}q_{3}{\rvert}\sim{\lvert}q_{0}{\rvert}\sim{\lvert}(\xi_{1},q_{1}){\rvert},\ \text{with}\ {\lvert}{\lvert}q_{i}{\rvert}-{\lvert}q_{j}{\rvert}{\rvert}\ll{\lvert}(\xi_{1},q_{1}){\rvert}\ \forall i,j\in\{0,1,2,3\},$
and this in turn implies that the $q_{i}$ cannot all have the same sign. Without loss of generality, let us assume that we have the following sign configuration: $(q_{1},q_{2},q_{3},q_{0})=(+,-,+,+)$. Then the assumption ${\lvert}\xi_{1}+\xi_{3}{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$ allows us to infer the pointwise bound
${\lvert}\xi_{0}{\rvert}\langle(\xi_{0},q_{0})\rangle^{s}$ $\lesssim{\lvert}\xi_{1}+\xi_{3}{\rvert}^{\frac{1}{4}}\langle(\xi_{1},q_{1})\rangle^{s-}\langle(\xi_{3},q_{3})\rangle^{\frac{5}{12}+}{\lvert}\xi_{0}+(-\xi_{2}){\rvert}^{\frac{1}{4}}\langle(\xi_{0},q_{0})\rangle^{0-}$
$\ \ \ \cdot\langle(-\xi_{2},-q_{2})\rangle^{\frac{5}{12}+},$
and undoing Plancherel, followed by Hölder’s inequality and a subsequent application of (30) and (33), yields
$I_{f}$ $\lesssim{\lVert}I_{x}^{\frac{1}{4}}P^{1}(J^{s-}u_{1}J^{\frac{5}{12}+}u_{3}){\rVert}_{L_{txy}^{2}}{\lVert}I_{x}^{\frac{1}{4}}P^{1}(J^{0-}fJ^{\frac{5}{12}+}\tilde{u}_{2}){\rVert}_{L_{txy}^{2}}$
$\lesssim{\lVert}f{\rVert}_{X_{0,\frac{1}{2}-2\varepsilon}}{\lVert}u_{1}{\rVert}_{X_{s,\frac{1}{2}+\varepsilon}}{\lVert}u_{2}{\rVert}_{X_{\frac{5}{12}+,\frac{1}{2}+\varepsilon}}{\lVert}u_{3}{\rVert}_{X_{\frac{5}{12}+,\frac{1}{2}+\varepsilon}}$
with the last step being valid for every $s>\frac{5}{12}$, provided $\varepsilon>0$ is chosen sufficiently small. Here, the applicability of (30) and (33) is justified by the fact that, since ${\lvert}\xi_{1}+\xi_{3}{\rvert}\ll{\lvert}(\xi_{1},q_{1}){\rvert}$ and ${\lvert}q_{1}+q_{3}{\rvert}\sim{\lvert}(\xi_{1},q_{1}){\rvert}$, it follows that
${\lvert}3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}{\rvert}\sim{\lvert}(\xi_{1},q_{1}){\rvert}^{2}\gg{\lvert}\xi_{1}+\xi_{3}{\rvert}$.
It therefore remains to be examined what happens in the case ${\lvert}\xi_{1}+\xi_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$: Then it follows from (ii.3), from ${\lvert}\xi_{0}-\xi_{2}{\rvert}={\lvert}\xi_{1}+\xi_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$, and from the sign configuration of the $q_{i}$, that ${\lvert}q_{1}-q_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$, ${\lvert}q_{0}+q_{2}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$, and thus ${\lvert}q_{1}+q_{2}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$, which allows us to proceed exactly as in case (ii.3.2) (the active assumption (ii.3.3) allows the Airy $L_{tx}^{4}L_{y}^{2}$-estimate (19) to be applied beneficially four times here as well) to obtain the desired trilinear estimate for all $s>\frac{5}{12}$, provided $\varepsilon>0$ is chosen sufficiently small.
(ii.3.3.2) ${\lvert}\xi_{\text{max}}{\rvert}\sim{\lvert}(\xi_{1},q_{1}){\rvert}$:
(ii.3.3.2.1) ${\lvert}\xi_{i}+\xi_{j}{\rvert}\gtrsim{\lvert}\xi_{\text{max}}{\rvert}\text{, for all }(i,j)\in{\{(1,2),(1,3),(2,3)\}}$:
In this case, we have, due to the general assumption (ii.3),
@@
which allows us to control up to $\frac{3}{2}$- derivatives by means of the quantities $\langle\sigma_{i}\rangle$. Proceeding exactly as in case (ii.3.1), we thus arrive at (42), and this holds for every $s>\frac{1}{3}$, provided we choose $\varepsilon>0$ sufficiently small.
(ii.3.3.2.2) ${\lvert}\xi_{i}+\xi_{j}{\rvert}\gtrsim{\lvert}\xi_{\text{max}}{\rvert}\text{, for exactly two }(i,j)\in{\{(1,2),(1,3),(2,3)\}}$:
Without loss of generality, we may assume that the sign pattern is $(\xi_{1},\xi_{2},\xi_{3},\xi_{0})=(+,-,+,+)$ with ${\lvert}\xi_{1}+\xi_{2}{\rvert}\gtrsim{\lvert}\xi_{\text{max}}{\rvert}$ and ${\lvert}\xi_{2}+\xi_{3}{\rvert}\ll{\lvert}\xi_{\text{max}}{\rvert}$. If now ${\lvert}3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}{\rvert}\gtrsim{\lvert}\xi_{1}+\xi_{3}{\rvert}$ or ${\lvert}3(\xi_{1}+\xi_{2})^{2}-(q_{1}+q_{2})^{2}{\rvert}\gtrsim{\lvert}\xi_{1}+\xi_{2}{\rvert}$, then the bilinear refinement of the linear $L^{4}$-estimate applies exactly as in case (ii.3.1) and yields $s>\frac{1}{4}$. Thus we may restrict ourselves to the situation where
${\lvert}3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}{\rvert}\ll{\lvert}\xi_{1}+\xi_{3}{\rvert}$
@@
${\lvert}3(\xi_{1}+\xi_{2})^{2}-(q_{1}+q_{2})^{2}{\rvert}\ll{\lvert}\xi_{1}+\xi_{2}{\rvert}$
hold. Furthermore, we may assume that ${\lvert}\xi_{2}+\xi_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$, since otherwise, by (ii.3), we would have
${\lvert}R{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}{\lvert}\xi_{\text{max}}{\rvert}^{2},$
so that the resonance function - again as in case (ii.3.1) - already leads to at least $s>\frac{1}{3}$. In addition to that, we must have ${\lvert}q_{1}{\rvert}\sim{\lvert}(\xi_{1},q_{1}){\rvert}$ and ${\lvert}q_{3}{\rvert}\sim{\lvert}(\xi_{1},q_{1}){\rvert}$, and because of (ii.3) together with ${\lvert}\xi_{0}-\xi_{1}{\rvert}={\lvert}\xi_{2}+\xi_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$, we then also obtain ${\lvert}q_{2}{\rvert}\sim{\lvert}(\xi_{1},q_{1}){\rvert}$ and ${\lvert}q_{0}{\rvert}\sim{\lvert}(\xi_{1},q_{1}){\rvert}$ with ${\lvert}{\lvert}q_{1}{\rvert}-{\lvert}q_{0}{\rvert}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$ and ${\lvert}{\lvert}q_{2}{\rvert}-{\lvert}q_{3}{\rvert}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$. Otherwise (taking into account (ii.3) and the signs of the $\xi_{i}$) we would have
${\lvert}3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}{\rvert}\sim{\lvert}(\xi_{1},q_{1}){\rvert}^{2}\gg{\lvert}\xi_{1}+\xi_{3}{\rvert}$,
contradicting ${\lvert}3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}{\rvert}\ll{\lvert}\xi_{1}+\xi_{3}{\rvert}$. We are left with two cases:
  1. (1)
    If $q_{2}$ and $q_{3}$ have opposite signs, then ${\lvert}q_{2}+q_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$, and the smallness argument presented in case (ii.3.2) applies, yielding $s>\frac{5}{12}$ (the Airy $L_{tx}^{4}L_{y}^{2}$-estimate can be used four times due to the active assumption (ii.3.3)).
  2. (2)
    If $q_{2}$ and $q_{3}$ have the same sign, then, since $q_{1}+q_{2}+q_{3}=q_{0}$, the quantities $q_{0}$ and $q_{1}$ must have opposite signs. Combining ${\lvert}q_{2}-q_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$ with ${\lvert}q_{0}+q_{1}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$ then implies ${\lvert}q_{1}+q_{2}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$, so that we may again invoke the smallness argument to obtain $s>\frac{5}{12}$.
(ii.3.3.2.3) ${\lvert}\xi_{i}+\xi_{j}{\rvert}\gtrsim{\lvert}\xi_{\text{max}}{\rvert}\text{, for exactly one }(i,j)\in{\{(1,2),(1,3),(2,3)\}}$:
We again consider, without loss of generality, the sign configuration $(\xi_{1},\xi_{2},\xi_{3},\xi_{0})=(+,-,+,+)$. Since ${\lvert}\xi_{0}{\rvert}\lesssim{\lvert}\xi_{1}+\xi_{3}{\rvert}$, we may infer the pointwise bound
${\lvert}\xi_{0}{\rvert}\langle(\xi_{0},q_{0})\rangle^{s}$ $\lesssim{\lvert}\xi_{1}+\xi_{3}{\rvert}^{\frac{1}{24}}\langle(\xi_{1},q_{1})\rangle^{s-}\langle(\xi_{3},q_{3})\rangle^{\frac{11}{24}+}{\lvert}\xi_{0}+(-\xi_{2}){\rvert}^{\frac{1}{24}}\langle(\xi_{0},q_{0})\rangle^{0-}$
$\ \ \ \cdot\langle(-\xi_{2},-q_{2})\rangle^{\frac{11}{24}+},$
and under the assumption ${\lvert}3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}{\rvert}\gtrsim{\lvert}\xi_{1}+\xi_{3}{\rvert}^{\frac{1}{6}}$, undoing Plancherel, followed by Hölder’s inequality, and a subsequent application of (30) and (33) yields
$I_{f}$ $\lesssim{\lVert}I_{x}^{\frac{1}{24}}P^{\frac{1}{6}}(J^{s-}u_{1}J^{\frac{11}{24}+}u_{3}){\rVert}_{L_{txy}^{2}}{\lVert}I_{x}^{\frac{1}{24}}P^{\frac{1}{6}}(J^{0-}fJ^{\frac{11}{24}+}\tilde{u}_{2}){\rVert}_{L_{txy}^{2}}$
$\lesssim{\lVert}f{\rVert}_{X_{0,\frac{1}{2}-2\varepsilon}}{\lVert}u_{1}{\rVert}_{X_{s,\frac{1}{2}+\varepsilon}}{\lVert}u_{2}{\rVert}_{X_{\frac{11}{24}+,\frac{1}{2}+\varepsilon}}{\lVert}u_{3}{\rVert}_{X_{\frac{11}{24}+,\frac{1}{2}+\varepsilon}}$
with the last step requiring $s>\frac{11}{24}$ to hold. We may thus assume
${\lvert}3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}{\rvert}\ll{\lvert}\xi_{1}+\xi_{3}{\rvert}^{\frac{1}{6}}.$ (44)
From the general assumptions (ii.3), (ii.3.3), the signs of the $\xi_{i}$, ${\lvert}\xi_{1}+\xi_{2}{\rvert}\ll{\lvert}\xi_{0}{\rvert}$, and ${\lvert}\xi_{2}+\xi_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}$, it then follows that
${\lvert}q_{1}{\rvert}\sim{\lvert}q_{2}{\rvert}\sim{\lvert}q_{3}{\rvert}\sim{\lvert}q_{0}{\rvert}\sim{\lvert}(\xi_{1},q_{1}){\rvert}\ \text{with}\ {\lvert}{\lvert}q_{i}{\rvert}-{\lvert}q_{j}{\rvert}{\rvert}\ll{\lvert}\xi_{0}{\rvert}\ \forall i,j\in\{0,1,2,3\},$
and the constraint (44) additionally forces the sign patterns $(q_{1},q_{2},q_{3},q_{0})=\\ (\pm,\mp,\pm,\pm)$. If now
${\lvert}\xi_{1}+\xi_{2}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{5}{12}}\ \text{or}\ {\lvert}\xi_{2}+\xi_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{5}{12}},$
then by (ii.3), taking into account the signs of the $q_{i}$, we also obtain
${\lvert}q_{1}+q_{2}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{5}{12}}\ \text{or}\ {\lvert}q_{2}+q_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{5}{12}}.$
Passing to dyadic pieces (cf. the smallness argument in case (ii.3.1) and case (ii.3.2)) then yields
${\lVert}\partial_{x}(Q_{N_{1}}^{(\frac{5}{12})}(u_{1}u_{2})u_{3}){\rVert}_{X_{s,-\frac{1}{2}+2\varepsilon}}\lesssim N_{1}^{\frac{1}{2}+\frac{5}{12}-2s}\prod_{i=1}^{3}{\lVert}u_{i}{\rVert}_{X_{s,\frac{1}{2}+\varepsilon}}$
or
${\lVert}\partial_{x}(u_{1}Q_{N_{1}}^{(\frac{5}{12})}(u_{2}u_{3})){\rVert}_{X_{s,-\frac{1}{2}+2\varepsilon}}\lesssim N_{1}^{\frac{1}{2}+\frac{5}{12}-2s}\prod_{i=1}^{3}{\lVert}u_{i}{\rVert}_{X_{s,\frac{1}{2}+\varepsilon}},$
and after dyadic summation this gives the desired trilinear estimate (42), provided
$\frac{1}{2}+\frac{5}{12}-2s<0\Leftrightarrow\frac{11}{24}<s$
and $\varepsilon>0$ is sufficiently small. We may therefore assume
${\lvert}\xi_{1}+\xi_{2}{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{5}{12}}\ \text{and}\ {\lvert}\xi_{2}+\xi_{3}{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{5}{12}}.$
Moreover, the assumption ${\lvert}q_{1}-q_{3}{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{3}{4}}$ implies, by (ii.3), that ${\lvert}\xi_{1}-\xi_{3}{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{3}{4}}$, and therefore also
${\lvert}\xi_{1}+\xi_{2}{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{3}{4}}\ \text{or}\ {\lvert}\xi_{2}+\xi_{3}{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{3}{4}},$
as otherwise we would have
${\lvert}\xi_{0}{\rvert}^{\frac{3}{4}}\lesssim{\lvert}\xi_{1}-\xi_{3}{\rvert}\leq{\lvert}\xi_{1}+\xi_{2}{\rvert}+{\lvert}\xi_{2}+\xi_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{3}{4}}.$